Can you help me solve:

(-5x^2z^-3)^2/5xz^-2 x=-2 z=3

Answer by checkley71(8403)

(-5x^2z^-3)^2/5xz^-2
(-5*-2^2*3^-3)^2/5*-2*3^-2
(-5*4*1/27)^2/5*-2*1/9
(-20/27)^2/-10/9
(400/729)/-10/9
400/729*-9/10
-3600/7290=-360/729=-40/81 answer.